OpenStudy (anonymous):

How many times does the graph of the function below intersect or touch the x-axis? y=-3x^2+x+4

OpenStudy (anonymous):

when the graph touches x axis at that point y = 0

OpenStudy (anonymous):

so now when you put y =0 0 = -3x^2 + x +4 then you factor out

OpenStudy (anonymous):

\[-3x^2 + +4x - 3x +4=0\] \[-3x(x +1) +4(x +1)=0\] \[ (x+1) +(-3x +4) = 0 \] \[x = -1 , \frac{4}{3} \]

OpenStudy (anonymous):

this implies at two position \[when x = -1 { and} \frac{4}{3}\]

OpenStudy (anonymous):

Oh, I see. Thank you! I was just having trouble factoring after I realized that was what I was suppose to do.

OpenStudy (anonymous):

oh sorry in my second last reply i wrote \[(x + 1) + (-3x + 4)=0\] it's a typo instead it is \[(x +1) \ times (-3x + 4) = 0\]

OpenStudy (anonymous):

times means \[\times\]